OpenStudy (anonymous):

is anyone there for algebra 2

OpenStudy (anonymous):

just post question

OpenStudy (anonymous):

-log2x-3 i have to graph it and find the domain and range

OpenStudy (anonymous):

Need more detail, brackets. log of 2x, log (2x-3), log sub 2?

OpenStudy (anonymous):

the 2 is in the back of the log very small and the x is big

OpenStudy (anonymous):

\[y=\log_{2}x-3 \]

OpenStudy (anonymous):

Looks like that?

OpenStudy (anonymous):

Just plot a chart. log sub 2 of x=x to the power of something. So when x = 0, y=(2^0)-3

OpenStudy (amistre64):

2^y = x-3 (2^y) +3 = x is a good substitute..... right?